### v1

Abstract.
For $C^{1}$ diffeomorphisms with continuous invariant splitting without domination, we prove the existence of Pesin (un)stable manifold under the hyperbolicity of invariant measures.
Key words and phrases:
Stable manifold, non-uniform hyperbolicity, Lyapunov exponents, dominated splitting
2020 Mathematics Subject Classification:
37D05, 37D10, 37D25, 37D30
Y. Cao was partially supported by National Key R&D Program of China(2022YFA1005802) and NSFC 12371194; Z. Mi was partially supported by National Key R&D Program of China(2022YFA1007800) and NSFC 12271260; R. Zou was partially supported by National Key R&D Program of China(2022YFA1007800) and NSFC 12271386.
1. Introduction
The stable manifold theorem is one of the key tools in studying hyperbolic behavior on dynamical systems. It was firstly established in $C^{1}$ uniform hyperbolic systems (see e.g. [9, 6]), which asserts existence of stable/unstable manifolds tangent to the uniformly contracted/expanded sub-bundles, respectively. This has been developed in $C^{1+\alpha}(\alpha>0)$ non-uniformly hyperbolic systems by Pesin and others (see e.g. [8, 10, 15, 16, 19]), known as the Pesin’s stable manifold theorem. It is important to investigate if the $C^{1+\alpha}$ regularity assumption is essential or not.
Pugh [18] built a $C^{1}$ diffeomorphism which gives a counterexample to Pesin’s stable manifold theorem. More recently, Bonatti-Crovisier-Shinohara [5] showed that the non-existence of (un)stable manifolds is a generic phenomenon in the $C^{1}$ category. However, under the existence of domination on Oseledec’s splitting of the hyperbolic (without zero Lyapunov exponents) measure, one knows that the Pesin’s stable manifold holds true. This was announced by Mañé [11] in his ICM’s report, and was made precise by Abdenur-Bonatti-Crovisier[1, §8] and also by Avilar-Bochi [4, Theorem 4.7] with different approaches.
In the present paper, we remove the domination assumption in [1, 4], and show that the Pesin’s stable manifold theorem keeps valid for $C^{1}$ diffeomorphisms admitting invariant continuous splitting of hyperbolic measures.
Let $M$ be a $d$-dimensional compact Riemannian manifold without boundary. Denote by ${\rm Diff}^{1}(M)$ the space of $C^{1}$ diffeomorphisms endowed with the usual $C^{1}$-topology. Given $f\in{\rm Diff}^{1}(M)$ and an $f$-invariant Borel probability $\mu$. By Oseledec’s theorem [14], for $\mu$-almost every point $x\in M$, for every $v\in T_{x}M\setminus\{0\}$, there exists the limit
$\chi(x,v)=\lim_{n\to\pm\infty}\frac{1}{n}\log\|Df^{n}(x)v\|,$
called the Lyapunov exponents of $x$ at $v$.
A Pesin stable set at a point $x\in M$ is given by
$W^{s}(x)=\Big{\{}y\in M:\limsup_{n\to\infty}\frac{1}{n}\log d(f^{n}x,f^{n}y)<0\Big{\}}$
and analogously for the unstable set $W^{u}(x)$, replace $f$ by $f^{-1}$. We call $W^{s}(x)$/$W^{u}(x)$ the (global) Pesin stable/unstable manifold of $x$ if it is an injectively immersed sub-manifold. The Pesin’s stable manifold theorem asserts that for any hyperbolic invariant measure, for $\mu$-almost every point $x$, there exist Pesin stable manifold and Pesin unstable manifold corresponding to negative Lyapunov exponents and positive Lyapunov exponents with complementary dimensions.
Now we state our main result as follows. A more general version will be presented in Theorem C in Section 3.
Theorem A.
Let $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$. Assume that $\mu$ is an invariant measure whose Lyapunov exponents along $E$ are all positive and along $F$ are all negative.
Then, for $\mu$-almost every point $x$, there exists a Pesin stable manifold and a Pesin unstable manifold tangent to $E$ and $F$ respectively.
The Pesin’s stable manifold was carried out by Pliss [17] for $C^{1}$ diffeomorphisms when all Lyapunov exponents are all strictly negative, using the celebrated Pliss’s argument on hyperbolic times. It was then extended in [1, 4] to the case that the measure $\mu$ has positive and negative Lyapunov exponents simultaneously, whose Oseledec splitting $E\oplus F$ is uniformly dominated, i.e., there are $C>0$ and $\lambda\in(0,1)$ such that for $\mu$-almost every $x\in M$ and each $n\in{\mathbb{N}}$,
$\|Df^{-n}|_{E(f^{n}(x))}\|\cdot\|Df^{n}|_{F(x)}\|\leq C\lambda^{n}.$
Theorem A indicates that Pesin’s stable manifold theorem does hold for diffeomorphisms which are only $C^{1}$, as long as the $C^{1+\alpha}$ hypothesis is replaced by a continuous invariant splitting hypothesis on the measure’s Oseledec splitting. It extend the work of [1, 4] to $C^{1}$ diffeomorphims without domination.
When the measure has positive and negative Lyapunov exponents, different to [17], one needs the domination on measure’s Oseledec splitting to control the geometry of iterated disks tangent to stable/unstable direction. Both the original $C^{1+\alpha}$ hypothesis and the uniform domination hypothesis in [1, 4] provide such a domination. In particular, the main tools in [1, 4] are Plaque family theorem [9] and cone argument, which are guaranteed by uniform domination.
The main ingredient towards Theorem A is to create domination. A weak version of domination called averaged domination arises from the gap of Lyapunov exponents on sub-bundles. Though, weaker than uniform domination, it ensures that one can control the geometry well of iterated disks at hyperbolic times. This will be sufficient to obtain the unstable manifolds by considering the limit of these iterated disks.
2. Unstable manifolds from average domination
Let $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$. Denote by $d$ the distance on $M$ induced by its Riemannian structure. Assume that ${\rm dim}E=k$, let $d_{G}$ denote the distance in the Grassmannian manifold of $k$-dimensional subspaces in $TM$.
Given $x\in M$, a $C^{1}$-smooth embedded disk $W$ centered at $x$ is called a local stable manifold at $x$ if there exists $C>0$ and $\lambda\in(0,1)$ such that for any $n\geq 1$,
$d(f^{n}(y),f^{n}(z))\leq C\lambda^{n}d(x,y)$
whenever $y,z\in W$. Replacing $n$ to $-n$, one can define the local unstable manifold similarly.
Given $a\geq 0$, for every $x\in M$, define the $E$-direction cone of width $a$ at $x$ as follows:
$\mathscr{C}^{E}_{a}(x)=\left\{v=v^{E}\oplus v^{F}\in E(x)\oplus F(x):\|v_{F}\|\leq a\|v_{E}\|\right\}.$
We say a $C^{1}$ embedded sub-manifold $D$ is tangent to $\mathscr{C}^{E}_{a}$ if it has dimension ${\rm dim}E$ and $T_{x}D\subset\mathscr{C}^{E}_{a}(x)$ for every $x\in D$. In particular, we will say $D$ is tangent to $E$ everywhere if it is tangent to $\mathscr{C}^{E}_{0}$.
Denote by $m(A)=\inf_{\|v\|=1}\|Av\|=\|A^{-1}\|^{-1}$ the mini-norm of linear isomorphism $A$.
Given $\lambda>1$ and $n\in{\mathbb{N}}$, denote
[…]
When $x\in\Delta_{n}(\lambda,f)$, we say that $(x,f^{n}(x))$ is $\lambda$-average dominated. We will also study the set
$\Delta(\lambda,f)=\bigcap_{n\geq 1}\Delta_{n}(\lambda,f).$
Definition 2.1.
For any $\lambda_{1}>1,\lambda_{2}>1$, denote by $\mathcal{HD}(x,\lambda_{1},\lambda_{2},f)$ the set of times $n$ such that $x\in\Delta_{n}(\lambda_{2},f)$ and $n$ is a $\lambda_{1}$-hyperbolic time for $x$, i.e.,
$\prod_{j=n-k}^{n-1}m\left(Df|_{E(f^{j}x)}\right)\geq\lambda_{1}^{k},\quad 1\leq k\leq n.$
We write also $\mathcal{HD}(x,\lambda_{1},\lambda_{2})$ when there is no ambiguity.
By applying the average domination together with the backward contraction property on hyperbolic times, one can conclude the following result. See [13, Lemma 3.7, Proposition 3.8] for the proof.
Lemma 2.1.
Let $f$ be a $C^{1}$ diffeomorphism with a continuous invariant splitting $TM=E\oplus F$.
For any $\lambda_{1}>\sigma_{1}>1,\lambda_{2}>\sigma_{2}>1$, there exists $a>0$ and $r>0$ such that if $n\in\mathcal{HD}(x,\lambda_{1},\lambda_{2})$ for some $x\in M$, then for any sub-manifold $D$ tangent to $\mathscr{C}^{E}_{a}$ with radius larger than $r$ around $x$, for every $y\in D$ satisfying $d(f^{n}(x),f^{n}(y))\leq r$ we have
  • $d(f^{k}(x),f^{k}(y))\leq\sigma_{1}^{-n+k}d(f^{n}(x),f^{n}(y))$ for any $0\leq k\leq n$;
  • $d_{G}(T_{f^{k}(y)}f^{k}(D),E(f^{k}(y)))\leq\sigma_{2}^{-k}\cdot a$ for any $0\leq k\leq n$.
For an invariant measure $\mu$, we say a point $x\in M$ is $\mu$-generic if
$\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}(x)}\to\mu,\quad\textrm{as}~{}n\to\infty.$
If $\mu$ is ergodic, by Birkhoff’s ergodic theorem [20, Theorem 1.14], we know that $\mu$-almost every point $x$ is $\mu$-generic. The next result provides a condition on some $\mu$-generic point for the ergodic measure that guarantees the existence of unstable manifolds almost everywhere.
Theorem B.
Let $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$.
Assume that $\mu$ is an ergodic measure for which there exist $\lambda_{1}>1,\lambda_{2}>1$ and a $\mu$-generic point $x$ such that
$\limsup_{n\to\infty}\frac{1}{n}\#\left(\mathcal{HD}(x,\lambda_{1},\lambda_{2})\cap\{0,\cdots,n-1\}\right)>0.$
then at $\mu$-almost every point $x$ there exists a Pesin unstable manifold $W^{E,u}(x)$ tangent to $E$.
Moreover, for every $\chi\in(1,\lambda_{1})$ and $\mu$-almost every $x\in M$, there is a local Pesin unstable manifold $W^{E,u}_{\delta(x)}(x)$ of size $\delta(x)$ contained in the global Pesin unstable manifold $W^{E,u}(x)$ such that for any $y\in W^{E,u}_{\delta(x)}(x)$ and $n\in{\mathbb{N}}$,
$d(f^{-n}(x),f^{-n}(y))\leq T(x)\chi^{-n}d(x,y),$
where $T(x)$ and $\delta(x)$ are Borel functions.
Proof.
Let $\lambda>1$ and take $x$ as $\mu$-generic point whose density of $\mathcal{HD}(x,\lambda_{1},\lambda_{2})$ is positive. Thus, there exists $\theta\in(0,1)$ and a subsequence $\{n_{k}\}_{k\in{\mathbb{N}}}$ such that
(2.1) $\frac{1}{n_{k}}\#\left(\mathcal{HD}(x,\lambda_{1},\lambda_{2})\cap\{0,\cdots,n_{k}-1\}\right)>\theta,\quad\forall\,k\in{\mathbb{N}}.$
Write for simplicity the measures
$\mu_{n_{k}}=\frac{1}{n_{k}}\sum_{j=0}^{n_{k}-1}\delta_{f^{j}(x)},\quad k\in{\mathbb{N}}.$
Hence, we have $\mu_{n_{k}}\to\mu$ as $k\to+\infty$ by definition. For each $j\geq 0$, consider
$H_{j}=\begin{cases}\{f^{j}(x)\},&j\in\mathcal{HD}(x,\lambda_{1},\lambda_{2});\\ ~{}~{}~{}~{}\emptyset,&~{}~{}{\rm otherwise}.\end{cases}$
Observe that
$\frac{1}{n_{k}}\sum_{j=0}^{n_{k}-1}\delta_{f^{j}(x)}(H_{j})=\frac{1}{n_{k}}\#\left(\mathcal{HD}(x,\lambda_{1},\lambda_{2})\cap\{0,\cdots n_{k}-1\}\right).$
According to (2.1), one gets
$\mu_{n_{k}}\left(\bigcup_{0\leq j\leq n_{k}-1}H_{j}\right)\geq\frac{1}{n_{k}}\sum_{j=0}^{n_{k}-1}\delta_{f^{j}(x)}(H_{j})>\theta$
for ever $k\in{\mathbb{N}}$. Consider $H_{\infty}$ as the limit set of $\{H_{j}\}_{j\geq 0}$, that is
$H_{\infty}=\bigcap_{k\geq 1}\overline{\bigcup_{j\geq k}H_{j}}.$
One can check by definition directly that
$\mu(H_{\infty})\geq\limsup_{k\to\infty}\mu_{n_{k}}\left(\bigcup_{0\leq j\leq n_{k}-1}H_{j}\right).$
Thus, we obtain $\mu(H_{\infty})\geq\theta.$ Now we will show that there exists a local unstable disk centered at any $w\in H_{\infty}$.
Consider any $\chi\in(1,\lambda_{1})$ and $\sigma\in(1,\lambda_{2})$. Let $a>0$, $r>0$ be the constants given by Lemma 2.1. Take $D_{0}$ as a $C^{1}$ disk tangent to $\mathscr{C}^{E}_{a}$ centered at $x$ with radius $r$. From Lemma 2.1, we know that if $n\in\mathcal{HD}(x,\lambda_{1},\lambda_{2})$, then
  1. (1)
    $f^{n}(D_{0})$ contains a ball $D_{n}$ of radius $r$ centered at $f^{n}(x)$.
  2. (2)
    $d(f^{-k}(y),f^{-k}(z))\leq\chi^{-k}d(y,z)$ for any $0\leq k\leq n$ and $y,z\in D_{n}$.
  3. (3)
    $d_{G}\left(T_{f^{k}(y)}f^{k-n}(D_{n}),E(f^{k}(y))\right)\leq\sigma^{-k}\cdot a$ for any $0\leq k\leq n$ and $y\in f^{-n}(D_{n})$.
We claim that:
Claim.
$\{D_{n}\}_{n\in\mathcal{HD}(x,\lambda)}$ is uniformly equicontinuous in the $C^{1}$-topology.
Proof of the Claim.
Let $\mathcal{HD}(x,\lambda_{1},\lambda_{2})=\{h_{1},h_{2},\cdots,h_{m},\cdots\}$. Fix any $\varepsilon>0$. Since $E$ is continuous, there is $\delta_{1}>0$ such that
(2.2) $d_{G}(E(y),E(z))<\varepsilon/3$
whenever $d(y,z)<\delta_{1}$. By the property (3), there exists $i(\varepsilon)\in{\mathbb{N}}$ such that for any $i\geq i(\varepsilon)$, one has that
(2.3) $d_{G}(T_{y}(D_{h_{i}}),E(y))\leq\sigma^{-h_{i}}\cdot a<\varepsilon/3,\quad\forall\,y\in D_{h_{i}}.$
By (2.2) and (2.3), we see that for every $i\geq i(\varepsilon)$ and $y,z\in D_{h_{i}}$ satisfying $d(y,z)<\delta_{1}$,
$d_{G}(T_{y}D_{h_{i}},T_{z}D_{h_{i}})$$\leq$ $d_{G}(T_{y}D_{h_{i}},E(y))+d_{G}(E(y),E(z))$
$+$ $d_{G}(E(z),T_{z}D_{h_{i}})$
$<$ $\varepsilon.$
On the other hand, as every $D_{h_{i}}$ is $C^{1}$, there is $\delta_{2}>0$ such that for every $1\leq i\leq i(\varepsilon)$, if $y,z\in D_{h_{i}}$ with $d(y,z)<\delta_{2}$, then $d_{G}(T_{y}D_{h_{i}},T_{z}D_{h_{i}})<\varepsilon.$ Consequently, by taking $\delta=\min\{\delta_{1},\delta_{2}\}$, we get for every $i$ that
$d_{G}(T_{y}D_{h_{i}},T_{z}D_{h_{i}})<\varepsilon,$
whenever $d(y,z)<\delta$. This shows the desired result. ∎
For any $w\in H_{\infty}$, by construction there exists $\{h_{i}\}\subset\mathcal{HD}(x,\lambda_{1},\lambda_{2})$ such that $f^{h_{i}}(x)\to w$ as $i\to+\infty$. Since $\{D_{n}\}_{n\in\mathcal{HD}(x,\lambda_{1},\lambda_{2})}$ is uniformly equicontinuous by the claim above, according to the Ascoli-Arzela theorem, up to considering subsequences, there exists a $C^{1}$-disk $D^{u}(w,r)$ centered at $x$ of radius $r$ such that $D_{h_{i}}$ converges to $D^{u}(w,r)$ in $C^{1}$-topology. For any fixed $n\in{\mathbb{N}}$, choose $i_{0}\in{\mathbb{N}}$ large enough such that $h_{i}\geq n$ for every $i\geq i_{0}$. From the convergence of $D_{h_{i}}\to D^{u}(w,r)$, for any $y,z\in D^{u}(w,r)$, there exist $y_{h_{i}},z_{h_{i}}\in D_{h_{i}}$ such that $y_{h_{i}}\to y$, $z_{h_{i}}\to z$. By property (2), it follows that
$d\left(f^{-n}(y_{h_{i}}),f^{-n}(z_{h_{i}})\right)\leq\chi^{-n}d(y_{h_{i}},z_{h_{i}}),\quad\forall\,i\geq i_{0}.$
By passing to the limit as $i\to+\infty$, we get
$d(f^{-n}(y),f^{-n}(z))\leq\chi^{-n}d(y,z).$
From the arbitrariness of $n\in{\mathbb{N}}$ we know
$d(f^{-n}(y),f^{-n}(z))\leq\chi^{-n}d(y,z),\quad\forall\,n\in{\mathbb{N}}.$
In view of property (3) above, taking the limit as $i\to+\infty$ we obtain that $D^{u}(w,r)$ is tangent to $E$. Writing $W^{E,u}_{r}(w)=D^{u}(w,r)$ for every $w\in H_{\infty}$, which is a local Pesin unstable manifold tangent to $E$ of size $r$.
Set
$\Lambda_{0}=H_{\infty},\quad\Lambda_{k}=f^{k}(\Lambda_{0})\setminus\bigcup_{j=0}^{k-1}\Lambda_{j},\quad k\geq 1.$
Put $C=\min_{z\in M}m\left(Df|_{E(z)}\right)$, and then define
$\delta(x)=\sum_{k=0}^{\infty}\chi_{\Lambda_{k}}(x)C^{k}r,\quad T(x)=\sum_{k=0}^{\infty}\chi_{\Lambda_{k}}(x)\chi^{-k}C^{-k}.$
For each $x\in\Lambda_{k}$, since $f^{-k}(x)\in\Lambda_{0}=H_{\infty}$, which admits the local Pesin unstable manifold $W_{r}^{E,u}(f^{-k}(x))$ of size $r=\delta(f^{-k}(x))$. By a simple computation, we know $f^{k}(W_{r}^{E,u}(f^{-k}(x)))$ contains a ball of size $\delta(x)$ around $x$, denoted as $W^{E,u}_{\delta(x)}(x)$. Moreover, we have
$d(f^{-n}(y),f^{-n}(z))\leq T(x)\chi^{-n}d(y,z),\quad\forall\,n\in{\mathbb{N}},$
provided that $y,z\in W^{E,u}_{\delta(x)}(x)$. Since $\mu$ is ergodic, we have $\mu(\cup_{k\geq 0}\Lambda_{k})=1$. Therefore, we have shown that for $\mu$-almost every point $x\in M$, there exists a local Pesin unstable manifold $W^{E,u}_{\delta(x)}(x)$.
For $\mu$-almost every point $x\in M$, define
$W^{E,u}(x)=\bigcup_{n\geq 0}f^{n}\left(W^{E,u}_{\delta(f^{n}(x))}(f^{-n}(x))\right).$
Now we show it is a global Pesin unstable manifold at $x$. Observe first by construction that for every $y,z\in W^{E,u}(x)$, it holds that
$\limsup_{n\to+\infty}\frac{1}{n}\log d(f^{-n}(y),f^{-n}(z))\leq-\log\chi<0.$
Thus it suffices to show it is an injectively immersed submanifold. Since $\mu(\Lambda_{0})>0$, the ergodicity of $\mu$ implies that for $\mu$-almost every $x$, its backward iterates $\{f^{-n}(x)\}$ enters $\Lambda_{0}$ infinitely many times, whose local Pesin unstable manifolds have a radius $r$. Note also that for the size of the backward iterates of local Pesin unstable manifolds tends to zero for almost every point. Therefore, one can deduce that for any $m\in{\mathbb{N}}$ the union
$\bigcup_{0\leq k\leq m}f^{k}\left(W^{E,u}_{\delta(f^{k}(x))}(f^{-k}(x))\right)$
is contained in an embedded manifold $f^{n}(W^{E,u}_{r}(f^{-n}(x)))$ for some large $n$ within $f^{-n}(x)\in\Lambda_{0}$. This shows that $W^{E,u}(x)$ is an injectively immersed submanifold, which is a global Pesin unstable manifold at $x$. From the construction we know that both global and unstable Pesin unstable manifold are tangent to $E$.
3. Proof of Theorem A
Let $f\in{\rm Diff}^{1}(M)$. For a $Df$-invariant sub-bundle $E\subset TM$ and an invariant measure $\mu$, for $\mu$-almost every $x\in M$, denote the minimal and maximal Lyapunov exponents along $E$ by
$\chi_{E}^{-}(x)=\lim_{n\to+\infty}\frac{1}{n}\log m(Df^{n}|_{E(x)}),\quad\chi^{+}_{E}(x)=\lim_{n\to+\infty}\frac{1}{n}\log\|Df^{n}|_{E(x)}\|.$
Note that when $\mu$ is ergodic, $\chi_{E}^{-}(x)$ and $\chi_{E}^{+}(x)$ are constants at $\mu$-almost every $x\in M$, denoted them by $\chi_{E}^{-}(\mu)$ and $\chi_{E}^{+}(\mu)$, respectively.
In this section we will prove a stronger version of Theorem A stated below. According to the ergodic decomposition theorem [12, Theorem 6.4], it suffices to consider the ergodic case.
Theorem C.
Let $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$. If $\mu$ is an ergodic measure such that
$\chi_{E}^{-}(\mu)>\max\{0,~{}\chi_{F}^{+}(\mu)\},$
then at $\mu$-almost every point $x$ there exists a Pesin unstable manifold $W^{E,u}(x)$ tangent to $E$.
Moreover, for every $0<\varepsilon<\min\{\chi_{E}^{-}(\mu),\chi_{E}^{-}(\mu)-\chi_{F}^{+}(\mu)\}$ and $\mu$-almost every $x\in M$, there is a local Pesin unstable manifold $W^{E,u}_{\delta(x)}(x)$ of size $\delta(x)$ contained in the global Pesin unstable manifold $W^{E,u}(x)$ such that for any $y\in W^{E,u}_{\delta(x)}(x)$ and $n\in{\mathbb{N}}$,
$d(f^{-n}(x),f^{-n}(y))\leq T(x){\rm e}^{-\chi_{E}^{-}(\mu)n}{\rm e}^{\varepsilon n}d(x,y),$
where $\delta(x)$ and $T(x)$ are Borel functions.
By considering $f^{-1}$, using the same argument as in the proof of Theorem C, one can conclude the existence of Pesin’s stable manifolds tangent to $E$ when
$\chi_{E}^{+}(\mu)<\min\{0,~{}\chi_{F}^{-}(\mu)\}.$
The main goal of this section is to prove Theorem C, which is presented in Subsection 3.1.
3.1. Birth of averaged domination
Now we assume that $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$.
Given $\ell\in{\mathbb{N}}$, $\gamma_{1},\gamma_{2}\in{\mathbb{R}}$, denote
$\quad\Lambda_{\ell}(\gamma_{1},\gamma_{2},f)=\Big{\{}x:\frac{1}{n\ell}\sum_{i=0}^{n-1}\log m\left(Df^{\ell}|_{E(f^{i\ell}(x))}\right)\geq\gamma_{1},$
$\quad\quad\quad\quad\quad\quad\quad\quad~{}~{}\frac{1}{n\ell}\sum_{i=0}^{n-1}\log\|Df^{\ell}|_{F(f^{i\ell}(x))}\|\leq\gamma_{2},\quad\forall n\in{\mathbb{N}}\Big{\}}.$
By definition we have the next observation:
Lemma 3.1.
For every $\ell\in{\mathbb{N}}$, if $\gamma_{1}>\gamma_{2}$, then $\Lambda_{\ell}(\gamma_{1},\gamma_{2},f)\subset\Delta(\exp((\gamma_{1}-\gamma_{2})\ell),f^{\ell}).$
Lemma 3.2.
Given $\ell\in{\mathbb{N}}$ and $\gamma_{1}>\max\{0,\gamma_{2}\}$, then for every $\lambda\in(1,{\rm exp}(\ell\gamma_{1}))$, there exists $\theta\in(0,1)$ such that for every $x\in\Lambda_{\ell}(\gamma_{1},\gamma_{2},f)$ and for any $L\in{\mathbb{N}}$,
$\frac{1}{L}\#\left(\mathcal{HD}\left(x,\lambda,\exp((\gamma_{1}-\gamma_{2})\ell\right),f^{\ell})\cap\{0,\cdots,L-1\}\right)>\theta.$
Proof.
For every $L\in{\mathbb{N}}$, as
$\frac{1}{L}\sum_{j=0}^{L-1}\log m\left(Df^{\ell}|_{E(f^{j\ell}(x))}\right)\geq\ell\gamma_{1}$
by definition. For $\lambda\in(1,\exp(\ell\gamma_{1}))$, one can deduce from Pliss Lemma [2, Lemma 2.1] to get a density $\theta\in(0,1)$ such that for any $L\in{\mathbb{N}}$ there is a sequence of times $\{n_{i}\}_{i=1}^{j}\subset\{1,\cdots,L\}$ with $j\geq\theta L$ such that for every $1\leq i\leq j$,
$\prod_{j=n_{i}-k}^{n_{i}-1}m\left(Df^{\ell}|_{E(f^{j\ell}(x))}\right)\geq\lambda^{k},\quad 1\leq k\leq n_{i}.$
This together with Lemma 3.1 yields the desired result. ∎
Theorem 3.3.
Let $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$. If $\mu$ is an ergodic measure such that $\chi_{E}^{-}(\mu)>\chi_{F}^{+}(\mu)$, then
$\lim_{\ell\to\infty}\mu(\Lambda_{\ell}(\gamma_{1},\gamma_{2},f))=1$
whenever $\chi_{E}^{-}(\mu)>\gamma_{1}>\gamma_{2}>\chi_{F}^{+}(\mu)$.
As a consequence of Theorem 3.3 and Lemma 3.1, we have the next result, which suggests that by considering iterates, the average domination can birth from the gap of Lyapunov exponents between sub-bundles.
Corollary 3.4.
Let $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$. If $\mu$ is an ergodic measure such that $\chi_{E}^{-}(\mu)>\chi_{F}^{+}(\mu)$, then
$\lim_{\ell\to\infty}\mu(\Delta({\rm exp}((\gamma_{1}-\gamma_{2})\ell),f^{\ell}))=1$
whenever $\chi_{E}^{-}(\mu)>\gamma_{1}>\gamma_{2}>\chi_{F}^{+}(\mu)$.
Now we give the proof of Theorem C by applying Theorem B, Lemma 3.2 and Theorem 3.3.
Proof of Theorem C.
Take any $f$-ergodic measure $\mu$ such that
$\chi_{E}^{-}(\mu)>\max\{0,~{}\chi_{F}^{+}(\mu)\}.$
For any $0<\varepsilon<\min\{\chi_{E}^{-}(\mu),\chi_{E}^{-}(\mu)-\chi_{F}^{+}(\mu)\}$, take $\chi=\chi_{E}^{-}(\mu)-\varepsilon$. Then choose constants $\gamma_{1},\gamma_{2}$ satisfying
$\chi_{E}^{-}(\mu)>\gamma_{1}>\chi+\varepsilon/2>\chi>\gamma_{2}>\chi_{F}^{+}(\mu).$
It follows from Theorem 3.3 that
$\lim_{\ell\to\infty}\mu(\Lambda_{\ell}(\gamma_{1},\gamma_{2},f))=1.$
Therefore, one can fix $\ell\in{\mathbb{N}}$ so that $\mu(\Lambda_{\ell}(\gamma_{1},\gamma_{2},f))>0$.
From the choices of constants, we have $\gamma_{1}>\chi+\varepsilon/2>0$ and $\gamma_{1}>\gamma_{2}$. Thus, by applying Lemma 3.2, there is $\theta\in(0,1)$ so that for every $L>1$,
(3.4) $\frac{1}{L}\#\left(\mathcal{HD}(x,\lambda_{1},\lambda_{2},F)\cap\{0,\cdots,L-1\}\right)>\theta,$
where
$\lambda_{1}=\exp((\chi+\varepsilon/2)\ell),\quad\lambda_{2}=\exp((\gamma_{1}-\gamma_{2})\ell),\quad F=f^{\ell}.$
Note that $\mu$ maybe not $F$-ergodic. However, there exists some $F$-ergodic measure $\nu$ such that
$\mu=\frac{1}{\ell}\left(\nu+f_{\ast}\nu+\cdots f_{\ast}^{\ell-1}\nu\right).$
Since $\mu(\Lambda_{\ell}(\gamma_{1},\gamma_{2},f))>0$, up to considering another $f_{\ast}^{i}\nu$, we have
$\nu(\Lambda_{\ell}(\gamma_{1},\gamma_{2},f))>0.$
This implies that one can take a $\nu$-generic point $x\in\Lambda_{\ell}(\gamma_{1},\gamma_{2},f)$ w.r.t. $F$. Thus, together with (3.4), one knows that $(F,\nu,x)$ satisfies the assumption of Theorem B. Note also that
$1<\exp(\chi\ell)<\exp((\chi+\varepsilon/2)\ell)=\lambda_{1}.$
Hence, for $\mu$-almost every $x\in M$, there is a local Pesin unstable manifold $W^{E,u}_{\delta_{1}(x)}(x,F)\subset W^{E,u}(x,F)$ such that for any $y\in W^{E,u}_{\delta_{1}(x)}(x,F)$,
$d\left(F^{-n}(x),F^{-n}(y)\right)\leq T_{1}(x){\rm e}^{-\chi\ell n}d(x,y),\quad\forall\,n\in{\mathbb{N}},$
where $\delta_{1}(x)$ and $T_{1}(x)$ are Borel functions.
Going back to $f$, and using the ergodicity of $\mu$, one can find new Borel functions $\delta(x)$ and $T(x)$ so that for $\mu$-almost $x\in M$, there exists a local Pesin unstable manifold $W_{\delta(x)}^{E,u}(x,f)$ contained in its global Pesin unstable manifold $W^{E,u}(x,f)$, and for any $y\in W_{\delta(x)}^{E,u}(x,f)$,
$d(f^{-n}(x),f^{-n}(y))\leq T(x){\rm e}^{-\chi_{E}^{-}(\mu)n}{\rm e}^{\varepsilon n}d(x,y),~{}\forall\,n\in{\mathbb{N}}.$
We get the desired result. ∎
Now it remains to prove Theorem 3.3
3.2. Proof of Theorem 3.3
Instead, we prefer to prove the following result in terms of invariant measures.
Theorem 3.5.
Let $f\in{\rm Diff}^{1}(M)$ with a continuous invariant splitting $TM=E\oplus F$. If $\mu$ is an invariant measure for which there exist constants $\alpha,\beta$ satisfying $\chi_{E}^{-}(x)\geq\alpha>\beta\geq\chi_{F}^{+}(x)$ for $\mu$-almost every $x$, then
$\lim_{\ell\to\infty}\mu(\Lambda_{\ell}(\gamma_{1},\gamma_{2},f))=1$
whenever $\alpha>\gamma_{1}>\gamma_{2}>\beta$.
Since for the ergodic measure, its distinct Lyapunov exponents are constants almost everywhere, Theorem 3.3 can be deduced immediately from Theorem 3.5.
To proceed, we need an observation for sub-additive continuous sequence. Let $\Phi=\{\varphi_{n}\}_{n\in{\mathbb{N}}}$ be a sequence of continuous functions that is sub-additive, i.e., for any $x\in M$,
$\varphi_{n+m}(x)\leq\varphi_{n}(x)+\varphi_{m}(f^{n}(x)),\quad\forall\,n,m\in{\mathbb{N}}.$
Given $\gamma\in{\mathbb{R}}$ and $\ell\in{\mathbb{N}}$, put
$H_{\ell}(\Phi,\gamma)=\{x:\frac{1}{n\ell}\sum_{i=0}^{n-1}\varphi_{\ell}(f^{i\ell}(x))\leq\gamma,\quad\forall\,n\in{\mathbb{N}}\}.$
We have the following result.
Proposition 3.6.
Given $\gamma_{1}<\gamma_{2}$. Let $\Phi=\{\varphi_{n}\}_{n\in{\mathbb{N}}}$ be a sequence of sub-additive continuous functions. If $\mu$ is an invariant measure such that
$\lim_{n\to+\infty}\frac{\varphi_{n}(x)}{n}\leq\gamma_{1},\quad\mu\textrm{-a.e.}~{}x.$
Then
$\lim_{\ell\to\infty}\mu(H_{\ell}(\Phi,\gamma_{2}))=1.$
As a consequence of Proposition 3.6, we can prove Theorem 3.5 as follows.
Proof of Theorem 3.5.
Let us fix constants as in the assumption. For any $\ell\in{\mathbb{N}}$, consider the subset of $\Lambda_{\ell}(\gamma_{1},\gamma_{2},f)$ as follows
$\Lambda_{\ell}^{s}(\gamma_{2},f)=\{x:\frac{1}{n\ell}\sum_{i=0}^{n-1}\log\|Df^{\ell}|_{F(f^{i\ell}(x))}\|\leq\gamma_{2},\quad\forall\,n\in{\mathbb{N}}\}.$
Now we show
(3.5) $\mu(\Lambda_{\ell}^{s}(\gamma_{2},f))\to 1,\quad\textrm{as}~{}\ell\to\infty.$
Let
$\varphi_{n}(x)=\log\|Df^{\ell}|_{F(x)}\|,\quad n\in{\mathbb{N}},\quad x\in M.$
We see from chain’s rule that $\Phi=\{\varphi_{n}\}_{n\in{\mathbb{N}}}$ is a sub-additive continuous sequence. Since
$\lim_{n\to\infty}\frac{\varphi_{n}(x)}{n}=\chi_{F}^{+}(x)\leq\beta<\gamma_{2},$
one can apply Proposition 3.6 for $\beta$ and $\gamma_{2}$ to get that
$\mu(H_{\ell}(\Phi,\gamma_{2}))\to 1,\quad\textrm{as}~{}\ell\to\infty.$
Note that $\Lambda_{\ell}^{s}(\gamma_{2},f)=H_{\ell}(\Phi,\gamma_{2})$, we arrive the convergence (3.5).
Now take $\zeta_{n}(x)=-\log m(Df^{n}|_{E(x)})$ for every $x\in M$ and $n\in{\mathbb{N}}$. We know that $\{\zeta_{n}\}_{n\in{\mathbb{N}}}$ is also a sub-additive continuous sequence, which satisfies
$\lim_{n\to+\infty}\frac{\zeta_{n}(x)}{n}=-\lim_{n\to+\infty}\frac{1}{n}\log m\left(Df^{n}|_{E(x)}\right)=-\chi_{E}^{-}(x)\leq-\alpha<-\gamma_{1}.$
By Proposition 3.6 and meaning of $\zeta_{n},~{}n\in{\mathbb{N}}$, as above we obtain
(3.6) $\mu(\Lambda_{\ell}^{u}(\gamma_{1},f))\to 1,\quad\textrm{as}~{}\ell\to\infty,$
where
$\Lambda_{\ell}^{u}(\gamma_{1},f)=\{x:\frac{1}{n\ell}\sum_{i=0}^{n-1}\log m\left(Df^{\ell}|_{E(f^{i\ell}(x))}\right)\geq\gamma_{1},\quad\forall\,n\in{\mathbb{N}}\}.$
Combining (3.5) and (3.6) with the fact that $\Lambda_{\ell}(\gamma_{1},\gamma_{2},f)=\Lambda^{u}_{\ell}(\gamma_{1},f)\cap\Lambda_{\ell}^{s}(\gamma_{2},f)$, we obtain
$\lim_{\ell\to\infty}\mu(\Lambda_{\ell}(\gamma_{1},\gamma_{2},f))=1,$
which completes the proof.
To prove Proposition 3.6, we recall the next Pliss-Like Lemma given in [7, Lemma 5.8], which was firstly established in [3, Lemma A].
Lemma 3.7.
Given constants $\eta<\zeta<L$. Then for any $\varepsilon>0$, there exists $\rho=\rho(L,\eta,\zeta,\varepsilon)\in(0,1)$ such that for any sequence $\{a_{n}\}_{n\in{\mathbb{N}}}\subset{\mathbb{R}}$ with $a_{n}\leq L$ for every $n\in{\mathbb{N}}$, if
$\liminf_{n\to+\infty}\frac{1}{n}\#\{0\leq i\leq n-1:a_{i}<\eta\}>1-\rho,$
then there is a subset ${\mathcal{G}}\subset{\mathbb{N}}$ satisfying
$\limsup_{n\to+\infty}\frac{1}{n}\#\left({\mathcal{G}}\cap\{0,\cdots,n-1\}\right)>1-\varepsilon$
such that for any $j\in{\mathcal{G}}$, one has that
$\frac{1}{n}\sum_{i=0}^{n-1}a_{i+j}<\zeta,\quad\forall\,n\in{\mathbb{N}}.$
Now we give the proof of Proposition 3.6.
Proof of Proposition 3.6.
For any $\varepsilon>0$, take $\varepsilon^{\prime}>0$ so that $(1-\varepsilon^{\prime})^{2}>1-\varepsilon$. We will apply Lemma 3.7 by taking
$L=\max_{x\in M}\{\varphi_{1}(x)\},\quad\eta=\frac{1}{2}(\gamma_{1}+\gamma_{2}),\quad\zeta=\gamma_{2}.$
Let $\rho=\rho(L,\eta,\zeta,\varepsilon)\in(0,1)$ be the constant given by Lemma 3.7.
Define for each $\ell\in{\mathbb{N}}$ the set
$A_{\ell}=\{x:\frac{\varphi_{\ell}(x)}{\ell}<\eta\}.$
Since $\varphi_{n}(x)/n>\gamma_{1}>\gamma_{2}$ for $\mu$-almost every $x\in M$, there exists $\ell_{0}\in{\mathbb{N}}$ such that
(3.7) $\mu(A_{\ell})>1-\rho\varepsilon^{\prime},\quad\forall\,\ell\geq\ell_{0}.$
Now we fix any $\ell\geq\ell_{0}$. By Birkhoff’s ergodic theorem we know that the limit
$\psi(x):=\lim_{n\to+\infty}\frac{1}{n}\sum_{j=0}^{n-1}\chi_{A_{\ell}}\left(f^{j\ell}(x)\right)$
exists for $\mu$-almost every $x\in M$. Furthermore, together with (3.7) we have
(3.8) $\int\psi\,{\rm d}\mu=\mu(A_{\ell})>1-\rho\varepsilon^{\prime}.$
Consider $G=\{x\in K:1-\rho<\psi(x)\leq 1\}$, then applying (3.8) we obtain
$\mu(M\setminus G)=\mu(\{x:1-\psi(x)\geq\rho\})\leq\frac{1}{\rho}\int(1-\psi)\,{\rm d}\mu\leq\frac{\rho\varepsilon^{\prime}}{\rho}=\varepsilon^{\prime}.$
This shows that $\mu(G)>1-\varepsilon^{\prime}$. For any $x\in G$, put
$\mathcal{P}=\{i\geq 0:f^{i\ell}(x)\in A_{\ell}\}.$
Let
$a_{i}=\frac{\varphi_{\ell}(f^{i\ell}(x))}{\ell},\quad i\geq 0.$
By the sub-additivity, we have for every $i\in{\mathbb{N}}$ that
$a_{i}\leq\frac{1}{\ell}\sum_{j=0}^{\ell-1}\varphi_{1}(f^{i\ell+j}(x))\leq L.$
From the definition of $A_{\ell}$, we get $a_{i}<\eta$ for every $i\in\mathcal{P}$. Since $\psi$ is larger than $1-\rho$ on $G$, we know
$\lim_{n\to+\infty}\frac{1}{n}\#\left(\mathcal{P}\cap\{0,\cdots,n-1\}\right)>1-\rho.$
By applying Lemma 3.7, there exists ${\mathcal{G}}\subset{\mathbb{N}}$ satisfying
(3.9) $\limsup_{n\to+\infty}\frac{1}{n}\#\Big{(}{\mathcal{G}}\cap\{0,\cdots,n-1\}\Big{)}>1-\varepsilon^{\prime},$
and for any $j\in{\mathcal{G}}$, one has that
(3.10) $\frac{1}{n}\sum_{i=0}^{n-1}a_{i+j}<\zeta=\gamma_{2},\quad\forall\,n\in{\mathbb{N}}.$
Taking
$\tau(x,n):=\frac{1}{n}\sum_{j=0}^{n-1}\chi_{H_{\ell}(\Phi,\gamma_{2})}(f^{j\ell}(x)),\quad n\in{\mathbb{N}},$
by applying Birkhoff’s ergodic theorem, (3.9) and (3.10) suggest that $\lim_{n\to\infty}\tau(x,n)>1-\varepsilon^{\prime}$ for $\mu$-almost every $x\in G$. Since $\mu(G)>1-\varepsilon^{\prime}$, there exists $n_{0}\in{\mathbb{N}}$ such that the subset
$G_{0}=\left\{x\in G:\tau(x,n_{0})>1-\varepsilon^{\prime}\right\}$
satisfies $\mu(G_{0})>1-\varepsilon^{\prime}$. Now by the invariance of $\mu$, we get
$\mu(H_{\ell}(\Phi,\gamma_{2}))\geq\int_{G_{0}}\tau(x,n_{0}){\rm d}\mu\geq(1-\varepsilon^{\prime})\mu(G_{0})>(1-\varepsilon^{\prime})^{2}>1-\varepsilon.$
Up to now, we have shown $\liminf_{\ell\to\infty}\mu(H_{\ell}(\Phi,\gamma_{2}))\geq 1$, which implies
$\lim_{\ell\to\infty}\mu(H_{\ell}(\Phi,\gamma_{2}))=1.$
This completes the proof.